OpenStudy (anonymous):

can some one help on this please?? ill award and fan ! show whether Rolle's theorem can be applied to f(x) if it can determine any values of c in the interval [o,2 pi] for which f'(c)=0 f(x)=cos x

OpenStudy (anonymous):

whats f(x)

OpenStudy (anonymous):

you'll need to specify f(x) = ???

OpenStudy (anonymous):

oh sprry @Kanwar245 and @dpaInc

OpenStudy (anonymous):

f(x)=cos x

OpenStudy (anonymous):

sorry *

OpenStudy (anonymous):

do you remember the conditions needed to apply Rolle's Theorem?

OpenStudy (anonymous):

conditions: 1- f(x) has to be continuous on [0, 2pi] .... is it? 2- f(x) has to be differentiable on (0, 2pi) .... is it? 3- f(0) = f(2pi) .... is it?

OpenStudy (anonymous):

@dpaInc are you asking ??

OpenStudy (anonymous):

yes... i'm asking... if all those conditions are met, then u can use Rolle's Theorem.

OpenStudy (anonymous):

To satisfy Rolle's theroem, f(0)=f(2pi), f(x) is continuous in [0, 2pi], and is differentiable in (0, 2pi). So, to show whether Rolle's theorem can be applied to f(x)=cos(x), show cos(0)=cos(2pi). And cos(x) is continuous and differentiable...